# User Simulation for Evaluating Information Access Systems

### v1

@@ 7.4.1 Conversational Search
In a conversational search setting, the user receives a single answer, as opposed to a list of results presented in a SERP. Therefore, “it is likely that their next query will be dependent on what that answer is, whether it is to directly follow up with another query, to clarify what they wanted, to move to a different aspect of their need, or to stop the search altogether” [Lipani:2021:TOIS].
@@
Overall, this approach requires the following data components: (1) a sample of information needs (i.e., topics), (2) for each topic, a pre-defined set of subtopics, (3) subtopic-level relevance judgments, and (4) a dialogue dataset with subtopic annotations for the estimation of state transition probabilities.
Simulating Answers to Clarifying Questions
@@
One of the most studied forms of system initiative in conversational search is clarifying questions, which aim to elucidate the user’s information need [Aliannejadi:2019:SIGIR, Zamani:2020:WWW, Bi:2021:ICTIR].
@@
Salle:2021:ECIR simulate how a user would respond to clarifying questions that are in the form: “Are you looking for [facet]?” Their simulator, called CoSearcher, has two main components: a user intent model and a persona model. The user intent model represents the user’s information need and returns a boolean value depending on whether the clarifying question matches the user’s intent. This is implemented in [Salle:2021:ECIR] by fine-tuning a BERT model for binary classification. The persona model (referred to as user parameters in the original work) specifies personal user characteristics, namely cooperativeness and patience. Cooperativeness ($\in[0,1]$) represents the user’s willingness to help the system by giving an informative answer (e.g., “No, but you are close” or “No, I’m looking for [intent]”) as opposed to a minimal response (“Yes” or “No”). Patience specifies the maximum effort, measured in the number of turns, the user is willing to spend interacting with the system.
@@
CoSearcher targets a specific type of clarifying questions, which ask directly about a single facet, and does not consider conversational context. Further, it assumes that there is a predefined set of facets for each information need that the system can ask about. Sekulic:2022:WSDM address these limitations by fine-tuning a transformer-based large language model (LLM) for the task of answering clarifying questions. Specifically, they employ DoubleHead GPT-2 with language modeling and classification losses. The first part of the training input is given to the model as the sequence in[SEP]q[SEP]cq[bos]a[eos], where $in$ is the textual description of the user’s information need that is underlying the user’s query $q$, $cq$ is the clarifying question asked by the system, and $a$ is the answer given by the user; [bos] and [eos] are special tokens indicating the beginning and end of a sequence, respectively, and [SEP] is a separation token. The second part of the training input for the DoubleHead model is given another input sequence with a distractor answer, and a binary label indicating which of the answers is preferable. The distractor answers are sampled from the training dataset heuristically (e.g., if the answer starts with “Yes” then the distractor answer starts with “No,” and vice versa). At inference time, the above input sequence is given without the answer segment, which will be generated by the LLM. A conversation history-aware variant of this approach is also proposed, which is capable of generating answers to further follow-up clarification questions, by extending the input sequence given to the LLM with previous user and system utterances.
Most recently, Owoicho:2023:SIGIR extend upon the work in [Sekulic:2022:WSDM] and develop a user simulator that is capable of generating a variety of utterances: (1) queries to seek information, (2) answers to clarifying questions, and (3) feedback to system responses, by few-shot prompting a ChatGPT model. It is worth noting that while these LLM-based approaches generate answers that are fluent and natural-sounding, they work much like black boxes; the behaviour of the simulated user can be controlled only indirectly and only to a certain extent via training examples.
7.4.2 Conversational Recommendation
@@
The main task of conversational recommender systems is to elicit user preferences using natural language interactions, point users to potential items of interest, and process feedback by users on the made suggestions [Jannach:2021:CSUR]. Conversational recommender systems can naturally be framed in the classical sense of task-oriented dialogue systems: the goal is to find items that satisfy the set of constraints expressed by the user, which can be represented in terms of slot-value pairs: $C=\langle(slot_{1},value_{1}),\dots,(slot_{n},value_{n})\rangle$. These constraints are based on the user’s long-term and short-term preferences and may change depending on the context and the items that are available. Given the close resemblance to task-oriented dialogue systems, it is natural to discuss techniques along the main components of a modular simulator architecture depicted in Fig. 7.3.
@@
One of the main challenges from a user simulation perspective is to maintain consistency during the dialogue, both in terms of dialogue flow (i.e., collaborating with the system to progress towards the user goal) and preferences (i.e., not contradicting preferences that were disclosed earlier). Zhang:2020:KDD employ an agenda-based dialogue policy that is guided by an interaction model. The interaction model specifies (1) the set of user actions, (2) the allowed transitions between them, and (3) the expected system response (action) for each user action. The latter component allows the simulator to determine whether the system responds to the user with an appropriate action (i.e., “understood” the user); if not, then a replacement action can be pushed onto the agenda. The interaction model can be a general-purpose one, i.e., the QRFA model in Fig. 7.2, or created specifically for the problem of conversational recommendation, i.e., the CIR6 model in Fig. 7.1.
@@
The ensure the consistency of preferences, Zhang:2020:KDD introduce a preference model, which may be seen as part of the user modeling component in Fig. 7.3. The preference model maintains a knowledge structure with $(slot,value,pref)$ triples, which can be queried for preferences on a given slot (to answer questions like “What are your favourite movie genres?”) and on specific slot-value pairs (to answer questions like “Do you like action movies?”). Note that the slot can also be a specific type of entity with the value being the unique identifier of the entity (which allows for answering questions like “Have you seen Inception?” and “How did you like it?”). The preference model may be empty at the beginning of the conversation or seeded with initial preferences. Whenever it is queried for slots or slot-value pairs, either an existing value is returned, or a new value is generated and saved for future requests.
Specifically, Zhang:2020:KDD create a preference model grounded in actual user preferences, by randomly sampling a user, then subsampling item ratings of that user from a dataset of historical user-item interactions (the rest of the ratings are used as held-out data for automatic evaluation). To ensure the consistency of preferences, the notion of a personal knowledge graph (PKG) is utilized, which is a “resource of structured information about entities personally related to its user” [Balog:2019:ICTIR]. The PKG has two types of nodes: items and attributes. Preferences on items are taken from historical data; for attributes, preferences are inferred by aggregating the preferences of entities with that attribute.
@@
The simplest approach to NLG is to base it on templates. To increase language variety, for each intent a number of different articulations may be created [Zhang:2020:KDD].

### v2

@@ 7.5.1 Model-Based Approaches
In a conversational search setting, the user receives a single answer, as opposed to a list of results presented in a SERP. Therefore, “it is likely that their next query will be dependent on what that answer is, whether it is to directly follow up with another query, to clarify what they wanted, to move to a different aspect of their need, or to stop the search altogether” [Lipani:2021:TOIS]. This section focuses on model-based approaches for simulating such user queries.
@@
Note that this method is not concerned with the generation of natural language questions, but rather focuses on asking questions, from a predefined set, in a natural order. Overall, the following data components are required: (1) a sample of information needs (i.e., topics), (2) for each topic, a predefined set of subtopics, (3) user queries corresponding to those subtopics, (4) subtopic-level relevance judgments, and (5) a dialogue dataset with subtopic annotations for the estimation of state transition probabilities.
7.5.2 Data-Driven Approaches
@@
One of the most studied forms of system initiative in conversational search is clarifying questions, which aim to elucidate the user’s information need [Aliannejadi:2019:SIGIR, Zamani:2020:WWW, Bi:2021:ICTIR]. Data-driven approaches to simulating user responses to such questions leverage LLMs, either through specialized supervised training procedures or via in-context learning (i.e., prompting); see Table 7.3 for an overview.
Table 7.3: Overview of LLM-based user simulators for simulating responses to clarifying questions.
Reference Learning paradigm LLM Behavioural parameters
[Salle:2021:ECIR] Supervised fine-tuning BERT cooperativeness, patience
[Sekulic:2022:WSDM] Supervised fine-tuning GPT-2 -
[Owoicho:2023:SIGIR] In-context learning GPT-3 patience
[Wang:2024:WWW] Supervised fine-tuning T5 cooperativeness
CoSearcher [Salle:2021:ECIR]
@@
Salle:2021:ECIR simulate how a user would respond to clarifying questions that are in the form: “Are you looking for [facet]?” Their simulator, called CoSearcher, has two main components: a user intent model and a persona model. The user intent model represents the user’s information need and returns a boolean value depending on whether the clarifying question matches the user’s intent. This is implemented in [Salle:2021:ECIR] by fine-tuning a BERT model [Devlin:2019:NAACL] for binary classification. The persona model (referred to as user parameters in the original work) specifies personal user characteristics, namely cooperativeness and patience. Cooperativeness ($\in[0,1]$) represents the user’s willingness to help the system by giving an informative answer (e.g., “No, but you are close” or “No, I’m looking for [intent]”) as opposed to a minimal response (“Yes” or “No”). Patience specifies the maximum effort, measured in the number of turns, the user is willing to spend interacting with the system.
USi [Sekulic:2022:WSDM]
@@
CoSearcher targets a specific type of clarifying questions, which ask directly about a single facet, and does not consider conversational context. Further, it assumes that there is a predefined set of facets for each information need that the system can ask about. Sekulic:2022:WSDM address these limitations in their conversational user simulator, called USi, by fine-tuning a transformer-based LLM for the task of answering clarifying questions. Specifically, they employ DoubleHead GPT-2 [Radford:2019:misc] with language modeling and classification losses, and utilize existing datasets: Qulac [Aliannejadi:2019:SIGIR] and ClariQ [Aliannejadi:2020:arXiv]. The first part of the training input is given to the model as the sequence in[SEP]q[SEP]cq[bos]a[eos], where $in$ is the textual description of the user’s information need that is underlying the user’s query $q$, $cq$ is the clarifying question asked by the system, and $a$ is the answer given by the user; [bos] and [eos] are special tokens indicating the beginning and end of a sequence, respectively, and [SEP] is a separation token. The second part of the training input for the DoubleHead model is given another input sequence with a distractor answer and a binary label indicating which of the answers is preferable. The distractor answers are sampled from the training dataset heuristically (e.g., if the answer starts with “Yes” then the distractor answer starts with “No,” and vice versa). At inference time, the above input sequence is given without the answer segment, which will be generated by the LLM. A conversation history-aware variant of this approach is also proposed, which is capable of generating answers to further follow-up clarification questions, by extending the input sequence given to the LLM with previous user and system utterances.
ConvSim [Owoicho:2023:SIGIR]
Owoicho:2023:SIGIR extend upon the work of Sekulic:2022:WSDM and develop ConvSim, a user simulator that is capable of not only answering clarifying questions but also providing explicit feedback to system responses, by few-shot prompting a GPT-3 model [Brown:2020:NeurIPS]. All generated utterances need to be in line with the underlying information need $in$. Specifically, at a given conversational turn $t$, it is assumed that there exists a turn-level description of the information need $in_{t}$—this is created manually by expert annotators in [Owoicho:2023:SIGIR]. Then, the LLM is prompted with the respective task description (i.e., whether to answer a clarifying question or provide feedback), the turn-level information need $in_{t}$, sample transcripts with the desired behaviour, and the conversation history up to turn $t$. ConvSim is capable of simulating multi-turn interactions; following Salle:2021:ECIR, the approach considers patience, i.e., the maximum number of turns of feedback the simulated user is willing to provide.
Type+QA [Wang:2024:WWW]
Wang:2024:WWW present a user simulator by fine-tuning a pretrained T5 model [Raffel:2020:JMLR] for generating answers to clarifying questions, using the same datasets as Sekulic:2022:WSDM. Based on the analysis of low-scoring examples in the development set, they identify wrong answer types as the most common error, i.e., the model generates an answer with the wrong meaning, for example, “no” instead of “yes”. They propose a two-step model to address this issue. First, the answer type is predicted using a 4-way categorization (yes, no, open, irrelevant) by fine-tuning a pretrained text classification model (RoBERTa [Liu:2019:arXiv]). This prediction is turned into a decoding constraint at inference time: “if the predicted answer type is ‘yes’ or ‘no,’ then the generation starts with ‘yes’ or ‘no’ correspondingly” [Wang:2024:WWW]. Based on the realization that simulating user responses to clarifying questions is similar to the problem of QA, further gains are obtained by performing the fine-tuning for answer generation on top of a T5 checkpoint that has been trained on QA datasets [Khashabi:2020:EMNLP]. Simulators are trained with two different levels of cooperativeness, by partitioning the training data into two subsets, cooperative and uncooperative, based on response length.
Analysis of Simulated Responses
Multiple recent studies have performed an analysis of simulated user utterances to gain insights and understand their limitations. Sekulic:2024:TIST compare two representatives of LLM-based user simulators: USi [Sekulic:2022:WSDM], which follows a specific supervised training procedure and ConvSim [Owoicho:2023:SIGIR], which utilizes in-context learning (i.e., prompting). The comparison is done using both automatic natural language generation (NLG) metrics and human judges. While ConvSim performs worse than USi on automatic NLG metrics, this outcome is attributed to the known unreliability of such measures [Novikova:2017:EMNLP]. Conversely, in manual evaluation, ConvSim outperforms USi in both naturalness and usefulness, with the performance difference becoming more pronounced over multiple turns of interaction. Qualitative analysis of the responses reveals a greater frequency of short answers in case of USi, while the majority of the answers by ConvSim aim to clarify and refine the underlying information need. Additionally, it is worth noting that USi is based on GPT-2, while ConvSim is based on GPT-3. Consequently, some of the observed differences may be attributable to differences in the capabilities of the underlying LLMs.
Wang:2024:WWW compare open LLMs (Llama2 and Flan) with a commercial LLM (GPT-4) and find that open LLMs in a zero-shot setting do not sufficiently understand the user simulation task. GPT-4, on the other hand, can generate seemingly perfect answers even in situations where humans fail to understand the meaning of the clarifying question. This is likely not the desired human-like behaviour.
7.6 User Simulation for Conversational Recommendation
@@
The main task of conversational recommender systems is to elicit user preferences using natural language interactions, point users to potential items of interest, and process feedback by users on the made suggestions [Jannach:2021:CSUR]. Conversational recommender systems can naturally be framed in the traditional sense of task-oriented dialogue systems: the goal is to find items that satisfy the set of constraints expressed by the user, which can be represented in terms of slot-value pairs: $C=\langle(slot_{1},value_{1}),\dots,(slot_{n},value_{n})\rangle$. These constraints are based on the user’s long-term and short-term preferences and may change depending on the context and the items that are available.
Below, we discuss work on user simulation for conversational recommender systems, categorized into model-based vs. data-driven approaches.
7.6.1 Model-Based Approaches
Given the close resemblance to task-oriented dialogue systems, it is natural to discuss techniques along the main components of a modular simulator architecture depicted in Fig. 7.1.
@@ 7.6.1 Model-Based Approaches
One of the main challenges from a user simulation perspective is to maintain consistency during the dialogue, both in terms of dialogue flow (i.e., collaborating with the system to progress towards the user goal) and preferences (i.e., not contradicting preferences that were disclosed earlier).
Zhang:2018:CIKM assume a simple “system ask user respond” workflow, where the system keeps asking the user questions until it is confident in the results. It starts with an initial user request, prompting the system to search for candidate items. If the system lacks sufficient confidence in the results, it generates a question for the user and refines the results based on their answer (along with any previously collected question-answer pairs). The questioning repeats until the system is confident enough to present the final results to the user. Zhang:2018:CIKM utilize textual item reviews and transform each review into a question-answer sequence to simulate the conversation that resulted in the purchase of the given item. The initial request is constructed from the product category. The questions ask about specific product aspects, according to the order in which they are mentioned, and the simulated user answers these questions with the values mentioned in the review.
@@
Zhang:2020:KDD employ an agenda-based dialogue policy that is guided by an interaction model. The interaction model specifies (1) the set of user actions, (2) the allowed transitions between them, and (3) the expected system response (action) for each user action. The latter component allows the simulator to determine whether the system responds to the user with an appropriate action (i.e., “understood” the user); if not, then a replacement action can be pushed onto the agenda. The interaction model can be a general-purpose one, i.e., the QRFA model in Fig. 7.4, or created specifically for the problem of conversational recommendation, i.e., the CIR6 model in Fig. 7.3.
@@
The ensure the consistency of preferences, Zhang:2020:KDD introduce a preference model, which may be seen as part of the user modeling component in Fig. 7.1. The preference model maintains a knowledge structure with $(slot,value,pref)$ triples, which can be queried for preferences on a given slot (to answer questions like “What are your favourite movie genres?”) and on specific slot-value pairs (to answer questions like “Do you like action movies?”). Note that the slot can also be a specific type of entity with the value being the unique identifier of the entity (which allows for answering questions like “Have you seen Inception?” and “How did you like it?”). The preference model may be empty at the beginning of the conversation or seeded with initial preferences. Whenever it is queried for slots or slot-value pairs, either an existing value is returned, or a new value is generated and saved for future requests. To create a preference model grounded in actual user preferences, Zhang:2020:KDD randomly sample a user from a dataset of historical user-item interactions, then subsample item ratings of that user (the rest of the ratings are used as held-out data for automatic evaluation). To ensure the consistency of preferences, the notion of a personal knowledge graph (PKG) is utilized, which is a “resource of structured information about entities personally related to its user” [Balog:2019:ICTIR]. The PKG has two types of nodes: items and attributes. Preferences on items are taken from historical data; for attributes, preferences are inferred by aggregating the preferences of entities with that attribute.